# Generic motives and motivic cohomology of fields

### v1

@@ 1.1. Guide for reading
The conventions and main notation of this paper are gathered in the next Section 1.2. A more detailed summary of the the theories used in these notes are given in the first section. This concerns Voevodsky’s motives over a field, the known computation of motivic cohomology, and the homotopical theory of pro-object using $\infty$-categories. This section might serve as a reference section for the reader in the remaining of this text. Note we have also provided the reader with details of useful notions for the reading of the paper, in the Appendix: Section 6 on cycle modules is mainly due to Rost, and Section 7 recalls the basics of Bousfield localization.
Section 3, is concerned with the theory of generic motives, described previously. The definition follows that given in the PhD of the author [Dég02], but with the advantages of the $\infty$-categorical language. The importance of generic motives come from their close relation with Voevodsky’s homotopy $t$-structure. The crucial propery is 3.2.10. Secondly, generic motives can be compared with Beilinson’s notion of cohomological motives at the generic point, which were shown to be crucial for the motivic $t$-structure in [Bei02]. At the price444theoretically for duality reasons, practically for using resolution of singularities of inverting the characteristic exponent $e$ of the base field $k$ we get an anti-equivalence of categories in 3.2.2. Section 5.3 finally recalls the rich functoriality of generic motives, and the link with cycle modules (3.3.7).
In Section 4, we give several computations of generic motives, and of morphisms of generic motives: respectively, the case of the function field of rational curves (§4.1), arbitrary curves (§4.2), surfaces, and arbitrary higher dimension (§4.3). The attentive reader will have noticed a tentative conjecture to describe morphisms of generic motives at the end of Section 4.1, which allows to make more precise a question asked in the PhD of the author (see 4.1.6).
In the last part of these notes, Section 5, we focus on motivic cohomology of fields. We first consider the case of number fields: as explained previously, we show how to deduce from Borel’s computation of the rank of their K-groups the one of rational motivic cohomology, by a detailed study of the involved isomomorphisms with respect to appropriate $\lambda$-structures. We then deduce from the computation of particular generic motives, obtained in Section Section 4, some facts about motivic cohomology of fields, and the conjecture mentioned above. The section closes with some interesting cohomology theory that could be of interests.
@@ 1.2. Conventions
Category theory. We will work with $\infty$-categories, as defined in [Lur09, Lur17, Cis19]. By convention, all our $\infty$-categories will be stable presentable monoidal $\infty$-categories. We will often say functor for $\infty$-functors, when it does not lead to confusion. Given a stable $\infty$-category $\mathcal{C}$, we will denote by $\operatorname{Map}_{\mathcal{C}}$ the mapping space functor, seen in the $\infty$-category of spectra (see e.g. [Gal23, Constr. 5.50]). We let $\operatorname{Ho}\mathcal{C}$ be the associated homotopy category, with its canonical triangulated structure ([Lur17, Th. 1.1.2.13]). We simply denote by $\operatorname{Hom}_{\mathcal{C}}$ the abelian groups of morphism in $\operatorname{Ho}\mathcal{C}$, so that $\operatorname{Hom}_{\mathcal{C}}(X,Y)=\pi_{0}\operatorname{Map}_{\mathcal{C}}(X,Y)$, where $\pi_{0}$ is the 00-th stable homotopy group of the corresponding spectrum.
The use of $\infty$-category is justified by the consideration of pro-objects in homotopy categories (such as Voevodsky motives) which is central in this paper; see Section 2.3.
@@
We say monoidal for symmetric monoidal. We say rigid or strongly dualizable for an object in such a monoidal category which admits a strong dual.
Given an abelian group $A$, we put $A_{\mathbb{Q}}=A\otimes_{\mathbb{Z}}\mathbb{Q}$ and $A_{e}=A[1/e]$ for $e>0$. If $\mathcal{C}$ is a stable $\infty$-category, we let $\mathcal{C}_{\mathbb{Q}}$ (resp. $\mathcal{C}_{e}$) be the $\mathbb{Q}$-linearisation (resp. $e$-localisation) of $\mathcal{C}$.
Algebraic geometry. We will fix a base field $k$ of characteristic exponent $e$: $e=1$ (resp. $p$) if $k$ is a $\mathbb{Q}$-algebra (resp $\mathbb{F}_{p}$-algebra for a prime $p$). When using $\mathbb{Z}$-coefficients, the field $k$ is assumed to be perfect. [^5] This assumption is used in the theory of motivic complexes as recalled in Section Section 2.1. In order to use de Jong’s resolution of singularities, we sometime use $\mathbb{Z}[1/e]$-coefficients. Then $k$ can be arbitrary as the categories becomes invariants under purely inseparable extensions.
We say smooth for smooth of finite type.
@@
Given a scheme $X$, we let $X^{(n)}$ (resp. $X_{(n)}$) be the set of points of $X$ of codimension (resp. whose reduced closure has Krull dimension) $n$.
@@
A function field over $k$ will be a separable finitely generated extension $F$ of $k$. Without precision, a valuation $v$ on $E/k$ will be assumed to be discrete and with ring of integers $\mathcal{O}_{v}$ a $k$-algebra essentially of finite type. When one wants to be precise, one says that $v$ is geometric. We also simply says that $(E,v)$ is a valued function field.
@@
Motives. The theory and notation of motivic complexes is recalled in Section Section 2.1. We denote byt $\mathbbm{1}_{k}$ the unit for the $\otimes$-structure on motivic complexes, effective and stable, and use the special notation for twists:
@@ 1.3. Acknowledgement
Among those who greatly inspired me in the theory of motives, Jacob Murre shares the place with Sasha Beilinson and Vladimir Voevodsky. In a more practical way, I thank Simon Pepin Lehalleur, Jan Nagel, Marc Levine, Bruno Kahn, Raphaël Ruimy and Swann Tubach for discussions that helped me greatly while conceiving the mathematics behind this paper. Special thanks go to Olivier Taïbi without whom Borel’s regulator would still be inaccessible to me. He helped me figure out its crucial compatibility with $\lambda$-operations. I also dearly thank all the participants of the covid 2021 workgroup on Borel’s paper housed in the ENS Lyon matrix.
@@
2. Preliminary recall
2.1. Notation from motivic homotopy
We will use Voevodsky’s theory of motivic complexes. This theory has been presented in many places: [VSF00, Chap. 5], [MVW06], [Dég07, Dég11]. We will use the notation of [Dég11, CD19] that we will now gather for the comfort of the reader.
@@ 2.1. Notation from motivic homotopy
The effective category. ([Dég11, §4.1, 5.1], [CD19, §11.1]) The category of effective motivic complexes $\operatorname{DM^{eff}}(k)$ is the $\mathbb{A}^{1}$-localization of the derived category of $\mathbb{Z}$-linear sheaves with transfers $\operatorname{Sh^{tr}}(k)$. As such, it can be identified with the full-subcategory of $\operatorname{\mathrm{D}}(\operatorname{Sh^{tr}}(k))$ made of $\mathbb{A}^{1}$-local complexes $K$. According to the fundamental theorem of Voevodsky, $K$ is $\mathbb{A}^{1}$-local if and only if its homotopy sheaves $\operatorname{\underline{H}}_{i}(K)$ are $\mathbb{A}^{1}$-invariant. One deduces that $\operatorname{DM^{eff}}(k)$ admits a $t$-structure, called the homotopy $t$-structure, whose heart $\operatorname{HI}(k)$ is made of $\mathbb{A}^{1}$-invariant sheaves with transfers. We use homological conventions for the homotopy $t$-structure.
The $\mathbb{A}^{1}$-localization functor of a complex $K$ in $\operatorname{\mathrm{D}}(\operatorname{Sh^{tr}}(k))$ is given by Suslin’s singular complex $\operatorname{C_{*}^{sus}}(K)$. The motive (aka motivic complex) of a smooth $k$-scheme $X$ of finite type is the complex $M(X)=\operatorname{C_{*}^{sus}}(\operatorname{\mathbb{Z}^{tr}}(X))$ where $\operatorname{\mathbb{Z}^{tr}}(X)$ is the sheaf with transfers represented by $X$ [^6] The definition of $\operatorname{M}(X)$ makes sense for a not necessarily smooth $k$-scheme $X$. In the non smooth case, $\operatorname{M}(X)$ has good properties only after inverting $e$. In these notes, we will only use $\operatorname{M}(X)$ for $X/k$ smooth. The tensor product on $\operatorname{DM^{eff}}(k)$ is characterized by the property $M(X)\otimes M(Y)=M(X\times Y)$ for $X$, $Y$ smooth. We denote by $\mathbbm{1}_{k}=M(k)$ the unit object for this tensor structure.
@@
The category of effective geometric motives $\operatorname{DM^{eff}_{gm}}(k)$ is the subcategory of $\operatorname{DM^{eff}}(k)$ made by finite sums, extensions, shifts and direct summands of motives of the form $M(X)$ for $X$ smooth. It coincides with the subcategory of compact objects of $\operatorname{DM^{eff}}(k)$.
2.1.2.
@@
There exists a universal (monoidal stable presentable) $\infty$-category $\operatorname{DM}(k)$ equipped with a functor $\Sigma^{\infty}:\operatorname{DM^{eff}}(k)\rightarrow\operatorname{DM}(k)$, called the infinite (Tate) suspension functor, such that $\Sigma^{\infty}\mathbbm{1}(1)$ is $\otimes$-invertible. [^7] See [Rob15, Def. 2.6]. The functor $\Sigma^{\infty}$ admits a right adjoint $\Omega^{\infty}$, the infinite (Tate) loop space functor. We will still simply denote by $\mathbbm{1}_{k}=\Sigma^{\infty}M(k)$ the unit for the tensor structure on $\operatorname{DM}(k)$.
There exists a unique $t$-structure on $\operatorname{DM}(k)$ such that $\Omega^{\infty}$ and the functor tensoring with $\mathbbm{1}\{1\}$ are $t$-exact. Its heart is the abelian category $\operatorname{HM}(k)$ of homotopy modules with transfers. Recall that a homotopy module with transfers is a $\mathbb{Z}$-graded sheaf with transfers $F_{*}=(F_{n})_{n\in\mathbb{Z}}$ equiped with a sequence of isomorphisms $(F_{n+1})_{-1}\simeq F_{n}$ for $n\in\mathbb{Z}$, where ?$_{-1}$ is Voevodsky’s $(-1)$-construction. [^8] This notion was introduced in the PhD of the author; see [Dég11, §2] for more details.
Voevodsky’s cancellation theorem ([Voe10]) asserts that $\Sigma^{\infty}$ is fully faithful. Therefore, we will use the same notation for $M(X)$ and $\Sigma^{\infty}M(X)$. One also deduces that $\Sigma^{\infty}$ induces a canonical fully faithful functor:
@@
As in the effective case, the category of geometric motives $\operatorname{DM_{gm}}(k)$ is the subcategory of $\operatorname{DM}(k)$ made by finite sums, extensions, shifts and direct summands of motives of the form $M(X)(n)$ for $X$ smooth, $n\in\mathbb{Z}$. It coincides with the subcategory of compact objects of $\operatorname{DM}(k)$.
2.1.3.
Duality and Chow motives. Let $X$ be a smooth $k$-scheme. We let $h(X):=M(X)^{\vee}=\operatorname{\underline{Hom}}(M(X),\mathbbm{1}_{k})$ be its canonical dual, the cohomological motive associated with $X$ [^9] In term of six functors, one has $h(X)=p_{*}(\mathbbm{1}_{k})$, $p:X\rightarrow\operatorname{Spec}(k)$ the structural map.
If $X/k$ is smooth projective of dimension $d$, it follows from [Dég08a, Th. 5.23] that $M(X)$ is rigid in $\operatorname{DM}(k)$ with dual $h(X)$. Moreover, there exists a canonical isomorphism $h(X)\simeq\operatorname{M}(X)(-d)[-2d]$. For a general smooth $k$-scheme $X$ of dimension $d$ (which can even be assumed to be non-smooth but of finite type), up to inverting the characteristic exponent $e$ of $k$, one deduces that $M(X)$ is rigid in $\operatorname{DM}(k)_{e}$ with dual $h(X)$. Moreover, one gets an isomorphism $h(X)\simeq M^{c}(X)(-d)[-2d]$, where $M^{c}(X)$ is the associated motive with compact support in $\operatorname{DM}(k)_{e}$ (see [VSF00, Chap. 5, 4.3.7] if $e=1$, and [CD15, Prop. 8.10, Rem. 7.4] in general). [^10] In terms of six operations, one has $M^{c}(X)=f_{*}f^{!}(\mathbbm{1}_{k})$, loc. cit. 8.10.
@@
$\operatorname{CHM}(k)^{op}\rightarrow\operatorname{Ho}\operatorname{DM_{gm}}(k)$

### v3

@@ 1.3. Guide for reading
The conventions and main notation of this paper are collected in Section 1.4. We use the language of $\infty$-categories, whose interest for generic motives has already been highlighted in the introduction. A comprehensive review of the foundational concepts required for the paper is provided in the Appendix, Section 6. In Section 2.1, we present Voevodsky’s theory of motives over a perfect field using the $\infty$-categorical framework. This section serves both as a reference for the remainder of the paper and as a concrete introduction to the $\infty$-categorical language through a significant case study. For readers already familiar with motivic complexes and willing to accept their $\infty$-categorical enhancement without further exposition, this section can be safely skipped on a first read.
Section 3 develops the theory of generic motives, extending the definitions given in [Dég02, Dég08b] using the enhanced framework of $\infty$-categories. We prove various properties of these objects, in close relation with Voevodsky’s homotopy $t$-structure. From a topos theoretic side, they can be considered as points according to 3.1.10. From the homotopical perspective, they behave like projective objects as explained in 3.1.14. We next give new presentations of the corresponding category (Section 3.2) and recall the link with Rost cycle modules (Section 3.3).
In Section 4, we set the stage with 4.2.7, addressing the structure of morphisms of generic motives. We then proceed to new computations of generic motives, extending the known case of rational curves (Section 4.2) to include arbitrary curves (Section 4.3) and surfaces (Section 4.4).
The final part of these notes, Section 5, focuses on motivic cohomology of fields. We begin with the case of number fields, showing in Section 5.1 how to derive the ranks of rational motivic cohomology groups from Borel’s computation of K-groups, with particular attention to the interplay with appropriately defined $\lambda$-structures. We then then apply the results on generic motives of purely transcendental extensions of rank one to prove B: see 5.2.2 and 5.2.4. The section concludes with a discussion of D (in the text: 5.2.6) and a brief exploration of a potentially fruitful cohomology theory that emerges naturally from these considerations.
@@ 1.4. Conventions
Category theory. We will work with $\infty$-categories, as defined in [Lur09, Lur17, Cis19]. As previously mentioned, a comprehensive account of the theory is provided in Section 6. As a rule, we use script letters for $\infty$-categories and roman letters for (ordinary) categories.
@@
Throughout, all our $\infty$-categories will be stable and presentable monoidal $\infty$-categories (see 6.3.5, 6.4.9). We will often refer to $\infty$-functors simply as functors when no confusion can arise. For a given $\infty$-category $\mathscr{C}$, we denote by $\operatorname{Map}_{\mathscr{C}}$ the mapping space functor (see 6.1.8 and 6.4.5). We let $\operatorname{Ho}\mathscr{C}$ be the associated homotopy category (see 6.1.6), equipped with its canonical triangulated structure (see 6.4.4). We write $\operatorname{Hom}_{\mathscr{C}}$ for the abelian groups of morphisms in $\operatorname{Ho}\mathscr{C}$, so that
$\operatorname{Hom}_{\mathscr{C}}(X,Y)=\pi_{0}\operatorname{Map}_{\mathscr{C}}(X,Y)$
where $\pi_{0}$ is the 00-th stable homotopy group of the corresponding spectrum. The use of $\infty$-categories is particularly justified by the consideration of pro-objects, for which we give a quick account in Section 6.6.
@@
With the exception of the appendix, we say monoidal for symmetric monoidal (applied to functors, categories, $\infty$-categories). We say rigid or strongly dualizable for an object in such a monoidal category that admits a strong dual.
We use homological conventions for $t$-structures (see also 6.4.10). Concretely, this means:
  • positive objects (denoted $M>0$) are stable under positive suspension,
  • $\operatorname{Hom}(M,N)=0$ for $M\geq 0$ and $N<0$,
  • one has a homotopy exact sequence: [^9] Recall the $\infty$-categorical terminology 6.4.3 $\tau_{\geq 0}M\rightarrow M\rightarrow\tau_{<0}M$ for all $M$.
Given an abelian group $A$, we put $A_{\mathbb{Q}}=A\otimes_{\mathbb{Z}}\mathbb{Q}$ and $A[1/e]=A\otimes_{\mathbb{Z}}\mathbb{Z}[1/e]$ for an integer $e>0$. If $\mathscr{C}$ is a stable $\infty$-category, we let $\mathscr{C}_{\mathbb{Q}}$ (resp. $\mathscr{C}[1/e]$) be its $\mathbb{Q}$-localization (resp. $e$-localization) — obtain by localizing with respect to morphisms $p.\mathrm{Id}$ for every prime $p$ (resp. morphisms $e.\mathrm{Id}$).
Algebraic geometry. We will fix a perfect base field $k$ [^10] This assumption can be dropped altogether if one works with $\mathbb{Z}[1/e]$-coefficients where $e$ is the characteristic exponent of $k$. This comes from the fact that the motivic categories then become invariant under inseparable extensions allowing a reduction to the perfect closure. See [Sus17].
We say smooth for smooth of finite type. We let $\mathrm{Sm}_{k}$ be the category of smooth $k$-schemes. We denote by $\mathrm{Sm}^{\mathrm{cor}}_{k}$ the additive category of smooth $k$-schemes with finite correspondences as morphisms (see below).
@@
Given a scheme $X$, we denote by $X^{(n)}$ (resp. $X_{(n)}$) the set of points of $X$ of codimension $n$ (resp. whose closure has Krull dimension $n$).
@@
A function field over $k$ will be a separable finitely generated extension $E$ of $k$. Unless otherwise specified, a valuation $v$ on $E/k$ is assumed to be discrete of rank $1$, with ring of integers $\mathcal{O}_{v}$ that is a $k$-algebra essentially of finite type. When one wants to be precise, one says that $v$ is geometric. We also simply say that $(E,v)$ is a valued function field.
@@
Motives. Our notation for motivic complexes is detailed in Section Section 2.1, together with the natural $\infty$-categorical enhancements. We denote by $\mathbbm{1}$ the unit for the $\otimes$-structure on motivic complexes, effective and stable, and use the special notation for twists:
@@ 1.5. Acknowledgement
Among those who greatly inspired me in the theory of motives, Jacob Murre shares a place with Sasha Beilinson and Vladimir Voevodsky.
I would like to sincerely thank Simon Pepin-Lehalleur, Jan Nagel, Marc Levine, Raphaël Ruimy, Robin Carlier, Swann Tubach, Tess Bouis, Jörg Wildeshaus and Rob de Jeu for various discussions and remarks which greatly contributed to the writing of this paper. Special thanks go to Olivier Taïbi without whom Borel’s regulator would still be inaccessible to me. He helped me figure out its crucial compatibility with $\lambda$-operations.
Finally, I am deeply grateful to the editors of this volume, Jan Nagel and Chris Peters, for their trust and unwavering patience, the opportunity to contribute, and their thorough help in refining the paper — from linguistic accuracy to the clarity of its mathematical formulations. I thank the anonymous referee for his vigilance.
@@
2. An $\infty$-categorical presentation of Voevodsky’s motivic theory
2.1. The $\infty$-category of motivic complexes and its stabilization
The purpose of this section is twofold.
First we give the notation and terminology from Voevodsky’s theory of motivic complexes over a perfect field $k$, following the assumptions laid out in Section 1.4. While the central definition can be given over an arbitrary base, as shown in [CD19, Part III], several key theorems — outlined below — require the base field $k$ to be perfect. [^11] As mentioned, the interested reader can find the necessary arguments to work over a non-perfect base field $k$ after inverting its characteristic exponent in [Sus17]. This has been presented in several sources: [VSF00, Chap. 5] (the original reference), [MVW06], [Dég07] and [BV08].
Second, we present the canonical $\infty$-categorical model of all the triangulated monoidal categories arising in Voevodsky’s theory, so as to take full advantage of the $\infty$-categorical framework. This is easy using the model categorical enhancements of the theory given in [Dég11, CD19], so that we use these works as references. We will also outline a way to get direct constructions using the theory of $\infty$-categories as recalled in Section 6.
In this presentation, we work with $\mathbb{Z}$-coefficients, but the framework can naturally be extended to $R$-linear coefficients for an arbitrary ring $R$. To indicate the coefficient rings, we use the notation $\operatorname{\mathscr{D}M^{eff}}(k,R)$ and $\operatorname{\mathscr{D}M}(k,R)$, with similar conventions for other $\infty$-categories.
@@ 2.1. The $\infty$-category of motivic complexes and its stabilization
The effective category. ([Dég07], [Dég11, §4.1, 5.1], [CD19, §11.1]) Recall that our base field $k$ is perfect. Voevodsky’s construction starts with the introduction of the additive category $\mathrm{Sm}^{\mathrm{cor}}_{k}$ of smooth $k$-schemes equipped with finite correspondences. This is a monoidal category whose tensor product is defined by the cartesian product on $k$-schemes and by the exterior product of finite correspondences. Using the graph of morphisms of smooth $k$-schemes, one obtains the so-called graph functor $\gamma:\mathrm{Sm}_{k}\rightarrow\mathrm{Sm}^{\mathrm{cor}}_{k}$.
The category of sheaves with transfers $\operatorname{Sh^{tr}}(k)$ over $k$ is defined as the category of abelian presheaves $F$ on $\mathrm{Sm}^{\mathrm{cor}}_{k}$ such that $F\circ\gamma$ is a Nisnevich sheaf. An important example is given by the sheaf with transfers $\operatorname{\mathbb{Z}^{tr}}(X)$, representable by a smooth $k$-scheme $X$. Its value over a smooth $k$-scheme $Y$ is the abelian group of finite correspondences $c(X,Y)$. For an arbitrary field, one shows that $\operatorname{Sh^{tr}}(k)$ is a Grothendieck abelian category. In particular, we can consider its derived $\infty$-category $\operatorname{\mathscr{D}}(\operatorname{Sh^{tr}}(k))$ (see 6.3.8). Note that, following our convention, this is indeed a presentable and stable $\infty$-category (see 6.4.6). Moreover, according to 6.4.11, it admits a canonical $t$-structure (see 6.4.10). We simply denotes by $\operatorname{\underline{H}}_{i}$ its $i$-th homology object.
One obtains the $\infty$-category of motivic complexes $\operatorname{\mathscr{D}M^{eff}}(k)$ by considering the $\mathbb{A}^{1}$-localization of $\operatorname{\mathscr{D}}(\operatorname{Sh^{tr}}(k))$: one formally inverts morphisms $\operatorname{\mathbb{Z}^{tr}}(\mathbb{A}^{1}_{X})\xrightarrow{p_{*}}\operatorname{\mathbb{Z}^{tr}}(X)$ induced by the canonical projection for an arbitrary smooth $k$-scheme $X$. It is obvious that its homotopy category is Voevodsky’s triangulated category $\operatorname{DM^{eff}}(k)$ of (unbounded) motivic complexes.
By the mechanism of localization in presentable $\infty$-categories (see 6.2.5), $\operatorname{\mathscr{D}M^{eff}}(k)$ can be identified with the full-$\infty$-subcategory of $\operatorname{\mathscr{D}}(\operatorname{Sh^{tr}}(k))$ made of motivic complexes $K$ that are $\mathbb{A}^{1}$-local: i.e. for any smooth $k$-scheme $X$, the induced morphism
$p^{*}:H^{*}_{\mathrm{Nis}}(X,K)\rightarrow H^{*}_{\mathrm{Nis}}(\mathbb{A}^{1}_{X},K)$
is an isomorphism. Let us recall the first key theorem of Voevodsky in this theory. This is where the assumption that $k$ is perfect occurs.
Theorem 2.1.2.
Consider the above notation and an arbitrary motivic complex $K$. Then $K$ is $\mathbb{A}^{1}$-local if and only if its homology sheaves $\operatorname{\underline{H}}_{i}(K)$ are $\mathbb{A}^{1}$-invariant.
Using the Nisnevich hypercohomology spectral sequence, one reduces to the case where $K$ is a sheaf with transfers in degree 00, in which case the proof can be found in [Dég07, Th. 4.5.1] or [Sus17] in the non perfect case. One deduces that the $\infty$-category $\operatorname{\mathscr{D}M^{eff}}(k)$ admits a $t$-structure, called the homotopy $t$-structure, whose heart $\operatorname{HI}(k)$ is made of $\mathbb{A}^{1}$-invariant sheaves with transfers, also called homotopy sheaves (see [Dég11, §5.1]).
Again by applying the general mechanism of localization in presentable $\infty$-categories, one gets the $\mathbb{A}^{1}$-localization functor denoted by $L_{\mathbb{A}^{1}}$, an exact endofunctor of $\operatorname{\mathscr{D}}(\operatorname{Sh^{tr}}(k))$. According to the previous theorem, $L_{\mathbb{A}^{1}}(K)$ can be identified with Suslin’s singular complex $\operatorname{C_{*}^{sus}}(K)$. The motive (aka motivic complex) of a smooth $k$-scheme $X$ of finite type is the complex $M(X)=\operatorname{C_{*}^{sus}}(\operatorname{\mathbb{Z}^{tr}}(X))$.
One gets a monoidal $\infty$-categorical structure on $\operatorname{\mathscr{D}M^{eff}}(k)$ by using the monoidal model structure of [CD09, Ex. 4.12] or [RØ08], and applying the construction of 6.5.12.121212We sketch below a procedure to avoid the recourse to one of these monoidal model categories. The tensor product on $\operatorname{\mathscr{D}M^{eff}}(k)$ is characterized by the fact that the $\infty$-functor $M:\operatorname{\mathscr{N}}\mathrm{Sm}^{\mathrm{cor}}_{k}\rightarrow\operatorname{\mathscr{D}M^{eff}}(k)$ is monoidal: $M(X)\otimes M(Y)=M(X\times Y)$.
@@
The $\infty$-category of effective geometric motives $\operatorname{DM^{eff}_{gm}}(k)$ is the sub-$\infty$-category of $\operatorname{DM^{eff}}(k)$ spanned by finite sums, extensions, shifts and direct summands of motivic complexes of the form $M(X)$ for $X$ smooth. Being geometric in this context is equivalent to being compact in the homotopy category $\operatorname{DM^{eff}}(k)$.
Remark 2.1.3.
If one seeks a direct $\infty$-categorical construction of the monoidal structure on $\operatorname{\mathscr{D}M^{eff}}(k)$, the following approach can be used:
  • Start with the presentable $\infty$-category $\operatorname{PSh}(\mathrm{Sm}^{\mathrm{cor}})$ of presheaves on the nerve of $\mathrm{Sm}^{\mathrm{cor}}$ as defined before 6.3.5.
  • Equip this $\infty$-category with a monoidal structure induced by the Day convolution product (6.5.7).
  • Localize with respect to the Nisnevich topology. [^13] Using Nisnevich distinguished squares is enough. One will still need to use [CD19, Prop. 10.3.9] to ensure that this localization is well-behaved.
  • Stabilize this monoidal $\infty$-category (see 6.5.9).
  • Perform a $\mathbb{Z}$-linearization procedure to get a presentable monoidal $\infty$-category whose underlying $\infty$-category is equivalent to $\operatorname{\mathscr{D}}(\operatorname{Sh^{tr}}(k))$.
After these steps, one can proceed with the $\mathbb{A}^{1}$-localization as described above, as it will be compatible with the monoidal structure (in the sense of 6.5.10).
Alternatively, one can directly use the derived $\infty$-category $\operatorname{\mathscr{D}}(\mathrm{Ab})$, equipped with its canonical monoidal structure. Indeed, one can start with the $\infty$-category of presheaves $\operatorname{\mathscr{F}un}\big{(}(\mathrm{Sm}^{\mathrm{cor}})^{\mathrm{op}},\operatorname{\mathscr{D}}(\mathrm{Ab})\big{)}$. As in the previous construction, it naturally inherits a monoidal structure via the Day convolution product (using [Gla16] in this context). One then proceeds as before, by applying Nisnevich localization and $\mathbb{A}^{1}$-localization. This approach bypasses the need for stabilization and $\mathbb{Z}$-linearization, and remains closer in spirit to the original construction.
2.1.4.
@@
Using the construction detailed in 6.5.8, there exists a universal (monoidal stable presentable) $\infty$-category $\operatorname{\mathscr{D}M}(k)$ along with a functor
$\Sigma^{\infty}:\operatorname{\mathscr{D}M^{eff}}(k)\rightarrow\operatorname{\mathscr{D}M}(k)$
called the infinite (Tate) suspension functor, such that $\Sigma^{\infty}\mathbbm{1}(1)$ is $\otimes$-invertible. The functor $\Sigma^{\infty}$ admits a right adjoint $\Omega^{\infty}$, the infinite (Tate) loop space functor. We will still simply denote by $\mathbbm{1}_{k}=\Sigma^{\infty}M(\operatorname{Spec}(k))$ the unit for the tensor structure on $\operatorname{\mathscr{D}M}(k)$. Objects of $\operatorname{\mathscr{D}M}(k)$ are called ordinary motivic spectra. Here, we add the adjective ordinary to make a difference with the objects of the stable homotopy $\infty$-category $\operatorname{\mathscr{SH}}(k)$ which are now frequently called motivic spectra. In this paper, there will be no risk of confusion so we just use the terminology motivic spectra for objects of $\operatorname{\mathscr{D}M}(k)$. We denote its homotopy category by $\operatorname{DM}(k)$, aligning with the notation used in the references indicated above.
There exists a unique $t$-structure on the $\infty$-category $\operatorname{\mathscr{D}M}(k)$ (see 6.4.10) such that the functors $\Omega^{\infty}$ and tensoring with $\mathbbm{1}\{1\}$ are $t$-exact (see [Dég12, Prop. 4.7]). Its heart is equivalent to the abelian category $\operatorname{HM}(k)$ of homotopy modules with transfers (we will simply call them homotopy modules in this work, as there is no risk of confusion with the objects of the heart of the stable homotopy category over $k$).
To describe the latter, we introduce Voevodsky’s $(-1)$-construction. Given a homotopy sheaf $F$, one defines a new homotopy sheaf $F_{-1}$, whose sections over a smooth $k$-scheme $X$ fit into the following split short exact sequence: [^14] Alternatively, one can show that $F_{-1}=\operatorname{\underline{Hom}}(\mathbb{G}_{m},F)$ using the internal Hom in $\operatorname{HI}(k)$.
(2.1.4.a) $0\rightarrow F(X)\xrightarrow{p^{*}}F(\mathbb{G}_{m,X})\rightarrow F_{-1}(X)\rightarrow 0$
One then defines a homotopy modules as a $\mathbb{Z}$-graded homotopy sheaf $F_{*}=(F_{n})_{n\in\mathbb{Z}}$ equipped with a sequence of isomorphisms $(F_{n+1})_{-1}\simeq F_{n}$ for $n\in\mathbb{Z}$. When $\mathbf{E}$ is a motivic spectrum, we denote by $\operatorname{\underline{H}}_{n,*}(\mathbf{E})$ its $n$-th homology object, viewed as a homotopy module, where the symbol * refers to the internal grading.
A cornerstone of the theory, Voevodsky’s Cancellation Theorem (see [Voe10a]) asserts that the functor $\Sigma^{\infty}$ is fully faithful. Consequently, we will use the same notation for both $M(X)$ and $\Sigma^{\infty}M(X)$. Moreover, it follows that $\Sigma^{\infty}$ induces a canonical fully faithful functor on the respective hearts:
@@
As in the effective case, the $\infty$-category of geometric motives $\operatorname{\mathscr{D}M_{gm}}(k)$ is the sub-$\infty$-category of $\operatorname{\mathscr{D}M}(k)$ spanned by finite sums, extensions, shifts and direct summands of motives of the form $M(X)(n)$ for $X$ smooth, $n\in\mathbb{Z}$. Again, being geometric is equivalent to being compact in the homotopy category $\operatorname{DM}(k)$.
2.1.5.
Duality and Chow motives. Let $X$ be a smooth $k$-scheme. We let $h(X):=M(X)^{\vee}=\operatorname{\underline{Hom}}(M(X),\mathbbm{1}_{k})$ be its canonical dual, the cohomological motive associated with $X$ [^15] In terms of the six functors formalism, one has $h(X)=p_{*}(\mathbbm{1}_{k})$, where $p:X\rightarrow\operatorname{Spec}(k)$ is the structural map.
If $X/k$ is smooth projective of dimension $d$, it follows from [Dég08a, Th. 5.23] that $M(X)$ is rigid in $\operatorname{\mathscr{D}M}(k)$ with dual $h(X)$. Moreover, there exists a canonical isomorphism $h(X)\simeq\operatorname{M}(X)(-d)[-2d]$. For a general smooth $k$-scheme $X$ of dimension $d$ (which can even be assumed to be non-smooth but of finite type), one also obtains that $M(X)$ is rigid in the $e$-localized category $\operatorname{DM}(k)[1/e]$, with dual $h(X)$. Moreover, one gets an isomorphism $h(X)\simeq M^{c}(X)(-d)[-2d]$, where $M^{c}(X)$ is the associated motive with compact support in $\operatorname{DM}(k)[1/e]$ (see [VSF00, Chap. 5, 4.3.7] if $e=1$, and [CD15, Prop. 8.10, Rem. 7.4] in general). [^16] In terms of six operations, one has $M^{c}(X)=f_{*}f^{!}(\mathbbm{1}_{k})$, loc. cit. 8.10.
@@
$\operatorname{CHM}(k)^{op}\longrightarrow\operatorname{DM_{gm}}(k)=\operatorname{Ho}\operatorname{\mathscr{D}M_{gm}}(k)$